OpenStudy (anonymous):

factor the trinomial completely ........25x^3-55x^2-60x

OpenStudy (anonymous):

Step one: take out the common factor (5). Step two: Use the ac method to finish factoring. Looking for a pair of factors of -60 that have a difference of -11 might lead you to -15 and 4.

OpenStudy (anonymous):

ok I get that part now.. I see how -15 and for would =11 when added and when multiplied it's 60

OpenStudy (anonymous):

-15 and 4 sorry

OpenStudy (anonymous):

\[25x^3-55x^2-60x=5x(5x^2-11x-12)\]\[5x(5x^2-15x+4x-12)=5x(5x(x-3)+4(x-3))\]\[=5x(5x+4)(x-3)\]

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

hold on let me check it out to see if I get it

OpenStudy (anonymous):

ok... I think I may understand it...I just really wish my text would download so I could try to read over it... I don't think you can make me understand anymore than I do

OpenStudy (anonymous):

thanks for trying to help me